# AlphaMaze: Enhancing Large Language Models’ Spatial Intelligence via GRPO

### v1

@@ Abstract
I Introduction
The ability to reason about visual information, particularly in spatial contexts, is a hallmark of intelligent systems. From navigating physical environments to interpreting complex diagrams, visual spatial reasoning is crucial for a wide range of tasks. While Large Language Models (LLMs) have achieved impressive performance in natural language processing and code generation, their capacity for genuine visual reasoning, especially spatial understanding and sequential decision-making in visual environments, remains a significant open question (Zhang et al., 2024; Ma et al., 2024). Current Vision-Language Models (VLMs) often excel at pattern recognition and object identification but may struggle with tasks requiring deeper spatial inference and step-by-step planning in visual domains (Ma et al., 2024). Bridging this gap and endowing standard LLMs with robust visual reasoning capabilities is a critical step towards more versatile and human-like AI.
@@ I Introduction
II Related Work
II-A Chain-of-Thought Reasoning in Language Models
Chain-of-Thought (CoT) prompting has emerged as a powerful technique to elicit complex reasoning from Large Language Models (Wei et al., 2022b). By prompting LLMs to ”think step by step,” CoT encourages the generation of intermediate reasoning steps, leading to improved performance on tasks requiring multi-step inference. Wei et al. (2022b) demonstrated that CoT prompting significantly enhances the ability of LLMs to solve arithmetic, commonsense reasoning, and symbolic reasoning tasks. Our work builds upon the concept of CoT reasoning, aiming to induce a similar step-by-step thought process in LLMs, but within the domain of visual spatial reasoning for maze navigation.
II-B Supervised Fine-Tuning for Visual and Spatial Tasks
@@
II-C Reinforcement Learning and GRPO for Reasoning and Reward Shaping
Reinforcement Learning from Human Feedback (RLHF) and its variants, such as Group Relative Policy Optimization (GRPO), have become increasingly important for aligning LLMs with human preferences and improving their reasoning abilities (Kwon et al., 2023).GRPO, as described by Shao et al. (2024) and implemented in DeepSeek-R1 (Guo et al., 2025), offers a computationally efficient approach to reinforcement learning by estimating advantages based on group scores, eliminating the need for a separate critic network. Reward function design is paramount in RLHF and GRPO, as it directly guides the model’s learning process. Carefully crafted reward functions can incentivize desired behaviors and shape the model’s policy towards optimal performance. Our work draws inspiration from the reward shaping strategies used in DeepSeek-R1 and adapts them to the context of visual maze navigation, designing reward components to encourage accuracy, valid movement sequences, and proper output formatting.
II-D DeepSeek-R1 and Emergent Reasoning through RL
@@
II-E Visual Reasoning and Maze Solving in AI
Maze solving has long been a benchmark task in Artificial Intelligence, serving as a testbed for various problem-solving and search algorithms (Janamian and Alam, 2023). Traditional approaches include graph search algorithms like Depth-First Search, Breadth-First Search, and A* (Lester, 2014-2024). More recently, AI techniques, particularly reinforcement learning and neural networks, have been applied to maze navigation (Zafrany, 2020). While prior work has explored maze solving using AI, our research focuses on a novel approach: teaching visual maze reasoning to standard language models through a tokenized visual representation and a combination of SFT and GRPO. This approach differs from traditional maze solvers by leveraging the inherent reasoning capabilities of LLMs and adapting them to process and reason about visual spatial information. Furthermore, research in neural-symbolic visual reasoning (Mao et al., 2023) explores combining neural networks with symbolic AI for visual tasks, offering a complementary perspective on integrating reasoning and visual processing.
III Methodology
III-A Tokenized Visual Maze Representation
@@
To enable the LLM to process maze information visually, we designed a tokenized input format that represents the maze grid, walls, origin, and target locations. Each cell in the maze is represented by a coordinate token <|row-col|>, e.g., <|0-0|> for the top-left cell. Wall information for each cell is encoded using tokens such as <|no_wall|>, <|up_wall|>, <|down_wall|>, <|left_wall|>, <|right_wall|>, <|up_down_wall|>, <|left_right_wall|>, <|up_left_wall|>, <|up_right_wall|>, <|down_left_wall|>, <|down_right_wall|>, <|up_down_left_wall|>, <|up_down_right_wall|>, <|up_left_right_wall|>, <|down_left_right_wall|>, <|up_down_left_right_wall|>. The origin and target locations are marked with <|origin|> and <|target|> tokens, respectively. Empty spaces within the maze representation are filled with <|blank|> tokens for consistent grid structure. This tokenization scheme provides a visual representation by explicitly encoding the spatial relationships between cells and the presence of walls, allowing the LLM to “see” the maze structure in a symbolic, tokenized form.
Example Maze Tokenization 1:
@@
III-B Baseline Models
To establish performance benchmarks for our approach, we employed three distinct baseline models, leveraging the DeepSeek-R1 (Guo et al., 2025) Distill-Qwen family of language models:
  • DeepSeek-R1-Distill-Qwen-7B: This represents a larger, 7 billion parameter model from the DeepSeek-R1 Distill-Qwen series, providing a strong performance reference point.
  • DeepSeek-R1-Distill-Qwen-1.5B: This is a smaller, 1.5 billion parameter counterpart, allowing us to assess the impact of model size within the same architecture family.
  • Direct Prediction Model (Customized DeepSeek-R1-Distill-Qwen-1.5B with SFT): Based on the 1.5B architecture, this baseline model was specifically fine-tuned using Supervised Fine-Tuning (SFT) to directly predict the entire sequence of movement tokens required to solve a maze in a single forward pass. The training objective was to minimize the cross-entropy loss between the predicted token sequence and the ground truth path. This direct prediction baseline serves to establish the performance of a standard LLM trained to output complete solutions without explicit step-by-step reasoning or reinforcement learning enhancement. The input to this baseline model was the tokenized maze representation, and the expected output was the concatenated sequence of movement tokens (e.g., <|up|> <|left|> <|right|> <|down|> ...).
@@ III-B Baseline Models
III-C Supervised Fine-Tuning (SFT) for Step-by-Step Reasoning
@@
For the SFT stage, we curated a training dataset. Mazes were synthetically generated with fixed sizes (5x5) and varied complexity level, controlled by parameters influencing path length and branching factor. Each maze in the dataset was paired with an annotated step-by-step solution, represented as a sequence of movement tokens (<|up|>, <|down|>, <|left|>, <|right|>). The Qwen 1.5B SFT model was then trained on this dataset. The training objective was to predict the next movement token at each step, conditioned on the maze input and the preceding movement tokens in the sequence as visually illustrated in Figure 2. This step-by-step prediction approach was designed to encourage the model to learn sequential reasoning for maze navigation.
III-D Group Relative Policy Optimization (GRPO) for Enhanced Reasoning
Following SFT, we applied Group Relative Policy Optimization (GRPO) to further enhance the model’s maze-solving capabilities and encourage more robust reasoning. The GRPO training utilized a smaller set of data than SFT state. We designed a reward function with the following components:
  • Correctness Reward (+0.2 per solution step): This reward is scaled according to the number of steps in the maze solution. Each valid movement step adds 0.2 points to the total score. For example, a solution requiring 4 steps earns a reward of $0.2\times 4=0.80.2 × 4 = 0.8$ points, incentivizing both accuracy and efficiency in navigation.
@@ III-D Group Relative Policy Optimization (GRPO) for Enhanced Reasoning
  • <think> Tag Reward (+0.25): This reward is given for correctly using the <think> tag in the output, ensuring completeness and consistency in the reasoning format.
@@
III-E Training Procedure and Pipeline
Our training pipeline consisted of two stages:
  1. 1.
    Supervised Fine-Tuning (SFT): The Qwen 1.5B SFT model was trained on the curated maze dataset for 10 epochs to learn step-by-step movement prediction for maze navigation.
  2. 2.
@@ III-E Training Procedure and Pipeline
    Group Relative Policy Optimization (GRPO): The SFT-trained model was subsequently further fine-tuned using the GRPO method, leveraging the designed reward function. Model checkpoints were saved at intervals of 200 steps.
@@
This two-stage pipeline mirrors the multi-stage training approach employed in DeepSeek-R1 (Guo et al., 2025), where initial RL training is followed by supervised fine-tuning for refinement. In our case, we used SFT before GRPO to provide a strong initial policy for the RL stage to build upon, focusing GRPO on refining reasoning and improving task-specific performance. All experiments were conducted using NVIDIA A6000 GPUs with LORA for parameter-efficient fine-tuning.
IV Experiments and Results
IV-A Dataset Details
The SFT training dataset comprises 500,000 synthetically generated mazes, created using the maze-dataset framework Ivanitskiy et al. (2023). This framework employs a randomized depth-first search algorithm, guaranteeing the existence of a solution path (connectivity) between the designated origin and target points within each maze. All mazes within the dataset are of a fixed size, 5x5 grids, providing a consistent spatial context for evaluation.
The full dataset is divided into three subsets: 500,000 mazes designated for SFT, providing the foundation for initial training; 16,000 mazes specifically allocated for GRPO, refining the model through policy optimization; and 30,000 mazes reserved for evaluation, ensuring a robust assessment of model performance.
IV-B MazeBench
@@
MazeBench is structured into three distinct difficulty levels – Easy, Medium, and Hard – to provide a graduated assessment of LLM performance. The difficulty is primarily determined by the approximate number of steps required for a viable solution path from the origin to the target. The benchmark comprises the following distribution:
  • Easy (50 mazes): Mazes in this category typically require a solution path of 1-4 steps. These mazes serve as a baseline assessment, testing fundamental pathfinding abilities.
  • Medium (40 mazes): These mazes necessitate solutions involving 5-8 steps, demanding more sophisticated planning and spatial reasoning.
  • Hard (10 mazes): Hard mazes present the most significant challenge, requiring solutions of 9-13 steps. These mazes test the limits of an LLM’s ability to handle long-range dependencies and complex spatial configurations.
As mentioned previously, the mazes are presented to the LLM in a tokenized input forma; the full details of this representation, including examples, are provided in Section III-A. The LLM is expected to produce output containing the following movement tokens: <|up|>, <|down|>, <|left|>, <|right|>. During evaluation, we will parse the LLM’s output to extract these tokens. The order of these tokens is crucial. The presence of extraneous characters, whitespace, or other tokens will not automatically invalidate the solution, provided that the correct sequence of movement tokens can be extracted. A solution is considered incorrect if the extracted sequence of movement tokens does not lead to the target or leads to an invalid state (e.g., attempting to move into a wall) is considered incorrect. The evaluation metric is the success rate: the percentage of mazes solved correctly.
IV-C Quantitative Results
IV-C1 Model Performance on MazeBench
As shown in Table I, the baseline model, trained for direct path prediction without explicit reasoning, achieved 0% accuracy on MazeBench. This highlights the necessity of step-by-step reasoning for the task. The SFT-only model reached a baseline of 86.0%, demonstrating the effectiveness of supervised fine-tuning for learning step-by-step maze navigation. Further enhancement with GRPO led to significant improvement, reaching 93.0% after 1600 steps of GRPO training.
TABLE I: Maze Solving Accuracy on MazeBench
Model SFT GRPO Score (%)

### v2

@@ Abstract
1   Introduction
The ability to reason about visual information, particularly in spatial contexts, is a hallmark of intelligent systems. From navigating physical environments to interpreting complex diagrams, visual spatial reasoning is crucial for a wide range of tasks. While Large Language Models (LLMs) have achieved impressive performance in natural language processing and code generation, their capacity for genuine visual reasoning, especially spatial understanding and sequential decision-making in visual environments, remains a significant open question (Zhang et al., 2024; Ma et al., 2024). Current Vision-Language Models (VLMs) often excel at pattern recognition and object identification but may struggle with tasks requiring deeper spatial inference and step-by-step planning
020040060080010001200140016001800200080808080858585859090909095959595GRPO stepsMazeBenchMazeBench scores with GRPO stepsMazeBenchTrendline (Linear Regression)±$\pm 1$ Std Dev
Figure 1: MazeBench scores over GRPO steps with a linear regression trendline and its ±$\pm 1$ standard deviation bounds.
in visual domains (Ma et al., 2024). Bridging this gap and endowing standard LLMs with robust visual reasoning capabilities is a critical step towards more versatile and human-like AI.
@@ 1   Introduction
2   Related Work
2.1   Chain-of-Thought Reasoning in Language Models
Chain-of-Thought (CoT) prompting has emerged as a powerful technique to elicit complex reasoning from Large Language Models (Wei et al., 2022b). By prompting LLMs to ”think step by step,” CoT encourages the generation of intermediate reasoning steps, leading to improved performance on tasks requiring multi-step inference. Prior research, including Wei et al. (2022b), Wei et al. (2023), Wang et al. (2023) prompting significantly enhances LLM performance on arithmetic, commonsense reasoning, and symbolic reasoning tasks. Our work builds upon the concept of CoT reasoning, aiming to induce a similar step-by-step thought process in LLMs, but within the domain of visual spatial reasoning for maze navigation.
2.2   Supervised Fine-Tuning for Visual and Spatial Tasks
@@
2.3   Reinforcement Learning and GRPO for Reasoning and Reward Shaping
Reinforcement Learning from Human Feedback (RLHF) and its variants have demonstrated significant efficacy in aligning Large Language Models (LLMs) with human preferences and enhancing their reasoning capabilities. However, RLHF faces substantial scalability challenges due to its resource-intensive nature and reliance on human feedback data. As an alternative approach, recent methodologies like Group Relative Policy Optimization (GRPO)(Kwon et al., 2023a) and Self-Play fIne-tuNing (SPIN) leverage self-play mechanisms, where models autonomously generate training signals and iteratively improve through self-competition Chen et al. (2024). These self-play approaches show promise in achieving human-level performance without the need for extensive human feedback, potentially offering a more scalable solution to the alignment challenge. GRPO, as described by Shao et al. (2024) and implemented in DeepSeek-R1 (Guo et al., 2025), offers a computationally efficient approach to reinforcement learning by estimating advantages based on group scores, eliminating the need for a separate critic network. Reward function design is paramount in RLHF and GRPO, as it directly guides the model’s learning process. Carefully crafted reward functions can incentivize desired behaviors and shape the model’s policy towards optimal performance. Our work draws inspiration from the reward shaping strategies used in DeepSeek-R1 and adapts them to the context of visual maze navigation, designing reward components to encourage accuracy, valid movement sequences, and proper output formatting.
2.4   DeepSeek-R1 and Emergent Reasoning through RL
@@
2.5   Visual Reasoning and Maze Solving in AI
Maze solving has long been a benchmark task in Artificial Intelligence, serving as a testbed for various problem-solving and search algorithms (Janamian and Alam, 2023). Traditional approaches include graph search algorithms like Depth-First Search, Breadth-First Search, and A* (Lester, 2014-2024). More recently, AI techniques, particularly reinforcement learning and neural networks, have been applied to maze navigation (Zafrany, 2020). While Chain-of-Thought (CoT) prompting has significantly enhanced complex reasoning capabilities in Large Language Models (LLMs) and Multimodal LLMs, it shows limitations in complex spatial reasoning tasks. Recent work by Microsoft introduces Multimodal Visualization-of-Thought (MVoT), which enables models to generate visual representations during their reasoning process, similar to human visual thinking (Li et al., 2025). This breakthrough demonstrates the potential of combining verbal and visual reasoning in AI systems.
Our research builds upon these advances, focusing on teaching visual maze reasoning to standard language models through a tokenized visual representation and a combination of SFT and GRPO. This approach differs from traditional maze solvers by leveraging the inherent reasoning capabilities of LLMs and adapting them to process and reason about visual spatial information. Furthermore, research in neural-symbolic visual reasoning (Mao et al., 2023) explores combining neural networks with symbolic AI for visual tasks, offering a complementary perspective on integrating reasoning and visual processing.
3   Methodology
3.1   Tokenized Visual Maze Representation
@@
To enable the LLM to process maze information visually, we designed a tokenized input format that represents the maze grid, walls, origin, and target locations. Each cell in the maze is represented by a coordinate token <|row-col|>, e.g., <|0-0|> for the top-left cell. Wall information for each cell is encoded using tokens such as <|no_wall|>, <|up_wall|>, <|up_down_wall|>, <|up_down_left_right_wall|>, .... The origin and target locations are marked with <|origin|> and <|target|> tokens, respectively. Empty spaces within the maze representation are filled with <|blank|> tokens for consistent grid structure. This tokenization scheme provides a visual representation by explicitly encoding the spatial relationships between cells and the presence of walls, allowing the LLM to “see” the maze structure in a symbolic, tokenized form.
Example Maze Tokenization 2:
@@
3.2   Baseline Models
To establish performance benchmarks for our approach, we employed three distinct baseline models, leveraging the DeepSeek-R1 (Guo et al., 2025) Distill-Qwen family of language models. We evaluate two distilled models: DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Qwen-1.5B. Additionally, a Direct Prediction baseline was established using a Supervised Fine-Tuning (SFT) approach on the DeepSeek-R1-Distill-Qwen-1.5B architecture. This model was trained to directly predict the complete sequence of movement tokens representing the solution path through a given maze. The training objective was the minimization of cross-entropy loss between the predicted token sequence and the ground truth solution. This baseline assesses the performance of a standard language model trained to generate complete solutions without intermediate reasoning steps or reinforcement learning techniques.
@@ 3.2   Baseline Models
3.3   Supervised Fine-Tuning (SFT) for Step-by-Step Reasoning
@@
For the SFT stage, we curated a training dataset. Mazes were synthetically generated with fixed sizes (5x5) and varied complexity level. The Qwen 1.5B SFT model was then trained on this dataset. The training objective was to predict the next movement token at each step, conditioned on the maze input and the preceding movement tokens in the sequence as visually illustrated in Figure 3. This step-by-step prediction approach was designed to encourage the model to learn sequential reasoning for maze navigation.
3.4   Group Relative Policy Optimization (GRPO) for Enhanced Reasoning
Following SFT, we applied Group Relative Policy Optimization (GRPO) to further enhance the model’s maze-solving capabilities and encourage more robust reasoning. The GRPO training utilized a smaller set of data than SFT state. We designed a reward function 3 components.
Correctness Reward (+0.2 per solution step): This reward is scaled according to the number of steps in the maze solution. Each valid movement step adds 0.2 points to the total score. For example, a solution requiring 4 steps earns a reward of $0.2\times 4=0.8$ points, incentivizing both accuracy and efficiency in navigation.
@@ 3.4   Group Relative Policy Optimization (GRPO) for Enhanced Reasoning
Thinking Reward (+0.25): This reward is given for correctly using the <think> tag in the output, ensuring completeness and consistency in the reasoning format.
@@
3.5   Training Procedure and Pipeline
Our training pipeline consisted of two stages. First, Supervised Fine-Tuning (SFT) was performed on the Qwen 1.5B model using a curated maze dataset for 10 epochs to learn step-by-step movement prediction for maze navigation. This phase established a strong initial policy, ensuring that the model could effectively interpret and respond to sequential movement tasks.
@@ 3.5   Training Procedure and Pipeline
Following SFT, Group Relative Policy Optimization (GRPO) was applied to refine the model’s performance. The SFT-trained model was further fine-tuned using LoRA Hu et al. (2021) with the GRPO method, implemented in Unsloth Daniel Han and team (2023) with VLLM Kwon et al. (2023b) for efficient inference. A carefully designed reward function guided the optimization process, and model checkpoints were saved every 200 steps to track improvements.
@@
This two-stage pipeline mirrors the multi-stage training approach employed in DeepSeek-R1 (Guo et al., 2025), where initial RL training is followed by supervised fine-tuning for refinement. In our case, SFT provided a robust starting point for reinforcement learning (RL), allowing GRPO to focus on refining reasoning capabilities and enhancing task-specific performance.
4   Experiments and Results
4.1   Dataset Details
The dataset is constructed through a multi-stage process involving generation, refinement, and augmentation. The process begins with the creation of a large initial pool of 530,000 synthetic mazes. These mazes are generated using the maze-dataset framework Ivanitskiy et al. (2023), which employs a randomized depth-first search algorithm. This algorithm ensures that every generated maze has a guaranteed solution path connecting the designated origin and target locations. Further details of algorithm used can be found at Appendix 1. All mazes within the dataset have a fixed size of 5x5 grids. From this extensive initial pool, a subset of 30,000 mazes is randomly selected and reserved as a held-out test set. This separation guarantees that the training and evaluation data are entirely distinct, preventing data leakage and enabling a robust assessment of model generalization.
The remaining 500,000 mazes form the basis for the various training datasets used in this work. This pool of mazes undergoes a multi-stage processing and augmentation procedure to create datasets specifically tailored for different training objectives.
Reset Dataset Creation: A significant portion of the training data focuses on teaching the model to recover from errors. To this end, a ”reset” dataset is created. This dataset is generated by algorithmically producing incorrect solution paths for the mazes. These incorrect paths are designed to be plausible but ultimately unsuccessful, either leading to dead ends or deviating from the correct solution. Importantly, they adhere to constraints: they do not revisit locations already visited within the incorrect attempt, and they avoid portions of the known correct solution path.
Associated with each incorrect path, a textual ”RESET” message is generated, simulating the feedback a system might provide upon encountering an error. The content of this message depends on whether the incorrect path terminates at a dead end (three surrounding walls) or simply deviates from the wrong route. These incorrect paths, along with their reset messages and the correct solution’s Chain-of-Thought (COT) reasoning, are combined. This process results in approximately 400,000 training examples where the model is presented with scenarios of initial failure(s) followed by a successful attempt after a ”reset.” The intent is to train the model to recognize incorrect trajectories and adapt its strategy. Illustrative examples of reset samples are provided in Appendix.
SFT Training Data Construction: The final Supervised Fine-Tuning (SFT) training dataset is a balanced combination of ”straight success” examples and ”retry” examples:
  • Straight Success Data (250,000 mazes): This portion consists of mazes where the model is expected to generate the correct solution path on the first attempt, without any resets.
  • Retry Data (250,000 mazes): This portion is drawn from the ”reset” dataset described above, providing examples where the model learns from incorrect attempts and subsequent resets. This combined 500,000-maze SFT set, encompassing success and error recovery, enables robust learning.
GRPO Training Data: The remaining 150,000 mazes from the original ”straight success” data pool are used to create a dataset for GRPO stage.
4.2   MazeBench
@@
MazeBench is a collection of 100 unique mazes, randomly selected from a larger test set containing 30,000 mazes. It is designed to evaluate the performance of large language models (LLMs) by categorizing mazes into different difficulty levels. Each maze requires the model to determine an optimal path from the starting point to the target, with difficulty primarily based on the number of steps needed to reach the goal.
The benchmark is structured into three levels: Easy, Medium, and Hard, ensuring a progressive assessment of an LLM’s pathfinding and problem-solving abilities. The components are described in Table 1.
Table 1: Maze Configuration by Difficulty Level
Category Number of Mazes Steps
Easy 50 1 – 4
Medium 40 5 – 8
Hard 10 9 – 13
Total 100 1 – 13
The Easy category consists of 50 mazes, each requiring between 1 and 4 steps to solve. These simpler mazes establish a baseline for evaluating fundamental navigation skills.
The Medium category includes 40 mazes that demand solution paths of 5 to 8 steps. These mazes introduce a higher level of complexity, requiring more advanced planning and spatial reasoning. Successfully solving them indicates an LLM’s ability to manage moderately intricate environments.
The Hard category comprises 10 mazes, each necessitating 9 to 13 steps to reach the target. These mazes present the greatest challenge, testing an LLM’s capacity to handle long-range dependencies and navigate complex spatial structures. Performance on this level reflects the model’s ability to process and reason over extended solution paths.
As mentioned previously, the mazes are presented to the LLM in a tokenized input format; the full details of this representation, including examples, are provided in Section 3.1. The LLM is expected to produce sequence of movement tokens. During evaluation, we will parse the LLM’s output to extract these tokens. The order of these tokens is crucial. The presence of extraneous characters, whitespace, or other tokens will not automatically invalidate the solution, provided that the correct sequence of movement tokens can be extracted. A solution is considered incorrect if the extracted sequence of movement tokens does not lead to the target or leads to an invalid state (e.g., attempting to move into a wall) is considered incorrect. The evaluation metric is the success rate: the percentage of mazes solved correctly.
4.3   Quantitative Results
4.3.1 Model Performance on MazeBench
As shown in Table 2, the baseline model, trained for direct path prediction without explicit reasoning, achieved 0% accuracy on MazeBench. This highlights the necessity of step-by-step reasoning for the task. The SFT-only model reached a baseline of 86.0%, demonstrating the effectiveness of supervised fine-tuning for learning step-by-step maze navigation. Further enhancement with GRPO led to significant improvement, reaching 93.0% after 1600 steps of GRPO training.
Table 2: Maze Solving Accuracy on MazeBench
Model SFT GRPO Score